OpenStudy (arbyj):

Someone me help solve this; -10x-9=-5x+8 (Please show work!!)

OpenStudy (anonymous):

So -10x-9=-5x+8 +5x +5x (you want to get the x's on one side of the equal sign) -5x-9=8 +9 +9 ( now you're getting the x by itself by adding 9 to both sides) -5x=17 (divide 17 by -5x) x=-3.4 or -17/5 or -3 2/5